

********************************************************************
UARS WEEKLY STATUS REPORT - 12 Nov 2004
********************************************************************
***********
SOLSTICE
***********
Nothing received.
***********
SUSIM
***********
SUSIM successfully gathered daily mid (1.1 nm) resolution, MgII,
occultation, solar pointing offset, and continuous monitoring
scans for time periods with UARS telemetry or tape recording.
***********
PEM
***********
HEPSA data files are version 2. Currently, HEPSA data between 1991274
and 2004199 has been reprocessed and archived at the GSFC DAAC.
MEPS data files are version 1. Processed MEPS data files at the
DAAC archive are between 1991274 and 2004199.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered.
The PEM data from the GUVI conjunction periods may be found at
http://wwwpem.space.swri.edu/guvi-pem.html. The additional electron data for
the October and November storms of 2003 has been moved and is now located at
http://climatology.space.swri.edu/uars_fromRudy.html. For those users of
SDDAS, update your UARS Meta data and these data times will appear
with full functionality. Proton data from these storms are now located at
http://climatology.space.swri.edu/UARS_images_from_Rudy_(Apr._2004).html
Requests have been received for storm spectral data. Overview plots were
produced and these plots may be found on the web at
http://climatology.space.swri.edu/UARS_images_postedJune2004.html
for the storms of October 29 of 1991, November 8-9 of 1991, April 6-7 of 2000,
March 31 - April 1 of 2001, April 11-12 of 2001, November 6-7 of 2001, and
November 11-12 of 2001.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 30 October 2004 (Orbit 71887, GMT Day 304) through
05 November 2004 (Orbit 71992, GMT Day 310).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 304/2349 71902 POWER OFF - ZEPS/NEPS
SUSIM 305/2204 71916 POWER OFF
SOLSTICE 305/2309 71916 FULLTIME TO DAY ONLY
SOLSTICE 309/1306 71971 DAY ONLY TO FULLTIME
SUSIM 309/1550 71972 POWER ON
A Reverse-to-Forward Yaw Around maneuver (number 134) was successfully
executed in Orbit 71941 beginning 307/ 15:08:42Z and completing at
307/ 15:45:52Z on 02 November 2004. All instruments were in their off
or standby mode (except for SOLSTICE in day-only). Due to the SA
position at 269 degrees, the effective night length of the spacecraft
after the maneuver was approximately 55.71 minutes. Post maneuver
analysis of data indicate that during the Yaw Around, the end-of-night
load bus voltage was 26.24 volts and the average depths of discharge
(DODs) on Batteries 2 and 3 were approximately 19.81 and 19.34 percent
respectively. The next (forward-to-reverse) YAW is scheduled for 02
December 2004 (DOY 337).
The beta angle ranged between the angles of 13.8 degrees to 12.3
degrees for this report period. The beta angle is now increasing
towards a maximum angle of 38 degrees on 18 November 2004 (DOY 323).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 0.64 to 0.96 degrees C. Additional
battery performance data for each battery on 29 October and 05
November 2004 are provided in Appendix D.
The clock error ranged from 10.0 to -10.6 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 03
PEM = 01
SUSIM = 02
SOLSTICE = 02
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 04
Three (3) Anomaly Reports remained OPEN at the end of this report
period (see Appendix E).
GROUND SYSTEM OPERATION
Four (4) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT took their first SMA pass with the spacecraft last week. Even
though the CMS is not able to receive the SMA contacts, the FOT was
able to schedule an SMA event on the UPS. The test was successful
with commands and dumps performing properly. Further use of the TDRS
SMA service will be postponed until the transfer issue between UPS and
CMS can be resolved.
The few test SMA events, which were scheduled via the workaround
feature on the UPS, were confirmed by scheduling this week. The test
events were deleted by the FOT, but the test verified the successful
execution of the workaround throughout the scheduling process.
Therefore no modifications to the FDF-provided UAV file will be
necessary.
MEETINGS
The FOT attended the biweekly meeting with the HRDI team on 03
November 2004 to discuss the progress in acquiring better attitude
knowledge for their data.
The follow-up meeting with FDF/Software Development scheduled for 04
November 2004 to discuss the rehost of FORMS 3 & 4 and the SMA issues
was cancelled.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 02 December 2004 (DOY 337).
This maneuver will be a Forward to Reverse Flight maneuver (Northern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 05 November 2004
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
71902 304/2349 04-310 PEM POWER OFF - ZEPS/NEPS
71916 305/2204 04-311 SUSIM POWER OFF
71916 305/2309 04-312 SOLSTICE SOLSTICE FULLTIME TO DAY ONLY
71941 307/1427 04-314 MACS YAW REV TO FWD
71942 307/1634 04-313 OBC/CLK CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
71962 309/0028 04-315 OBC/CLK CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
71971 309/1306 04-316 SOLSTICE SOLSTICE TO FULLTIME
71972 309/1550 04-317 SUSIM POWER ON : PART 1
71984 310/1155 04-317 SUSIM POWER ON
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 05 November 2004
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 05 November 2004
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
71942 307/1634 CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
71962 309/0032 CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 05 November 2004
29 October 2004 - GMT Day 303 Beta = 14.5 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.05/-4.31 +00.0/+00.0 07.0 0.00 0.0
2 05.6/- 39.2 +5.79/+5.14 +28.4/-13.6 26.2 1.03 20.3
3 -56.0/- 56.0 +4.50/+3.87 +27.2/-13.6 26.2 1.02 19.6
05 November 2004 - GMT Day 310 Beta = 12.3 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.05/-4.63 +00.0/+00.0 07.0 0.00 0.0
2 -22.4/- 72.8 +5.14/+4.50 +28.0/-12.4 26.6 1.02 18.9
3 00.0/- 78.4 +3.87/+2.91 +26.8/-12.0 26.4 1.03 18.3
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
-None this report period
APPENDIX E
UARS ANOMALY REPORTS
Week ending 05 November 2004
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-066 FOT 71891 304/0718 30832 Generic Late Acquisition
04-066 FOT 71955 308/1309 30832 Generic Late Acquisition
04-083 FOT 71977 310/0025 N/A UARS Eclipse Serve Failure
04-084 FOT 71985 310/1303 37501 TDRSS-275 dropouts
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-076 FOT 71783 297/0234 37341 Late Acquisition
04-077 FOT 71811 298/2220 37357 Multiple Data Dropouts
04-078 FOT 71826 299/2326 37358 DATA DROPOUTS
04-079 FOT 71834 300/1203 37396 TDRSS 275 Data Dropouts
04-080 FOT 71849 301/1200 37411 TDRSS 275 Dropouts
04-082 WSGT 71861 302/0647 37419 Drop Outs on the Return Link
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-066 FOT See List Oct 04 30832 Generic Late Acquisition
04-083 FOT 71977 310/0025 N/A UARS Eclipse Server Failure
04-084 FOT 71985 310/1303 37501 TDRSS-275 dropouts